Properties of a First-order Functional Lanuage with Sharing

Computation Structures Group Memo 347-1 15 December 1994 Supercedes CSG Memo 347

Zena M. Ariola University of Oregon

Arvind MIT

Funding for this work has been provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310 (MIT) and by the University of Oregon.